OpenStudy (hjutyu):

A ramp leads up to a building. The top of the ramp is 6 feet above the ground, and the bottom is 16 feet from the building as shown. What is the length of the ramp? Round your answer to the nearest tenth of a foot. A. 14.8 ft B. 17.1 ft C. 16.6 ft D. 22.0 ft

OpenStudy (hjutyu):

@Atrineas

OpenStudy (hjutyu):

@alphina

OpenStudy (hjutyu):

@Aamber

OpenStudy (hjutyu):

i need help please :(

OpenStudy (hjutyu):

@TheSmartOne

OpenStudy (hjutyu):

@Elsa213

TheSmartOne (thesmartone):

\(\bf\huge~~~~\color{#ff0000}{W}\color{#ff2000}{e}\color{#ff4000}{l}\color{#ff5f00}{c}\color{#ff7f00}{o}\color{#ffaa00}{m}\color{#ffd400}{e}~\color{#bfff00}{t}\color{#80ff00}{o}~\color{#00ff00}{O}\color{#00ff40}{p}\color{#00ff80}{e}\color{#00ffbf}{n}\color{#00ffff}{S}\color{#00aaff}{t}\color{#0055ff}{u}\color{#0000ff}{d}\color{#2300ff}{y}\color{#4600ff}{!}\color{#6800ff}{!}\color{#8b00ff}{!}\\\bf ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~Made~by~TheSmartOne\) Hey there!!! Since you are new here, read this legendary tutorial for new OpenStudiers!! http://openstudy.com/study#/updates/543de42fe4b0b3c6e146b5e8 @hjutyu @~ You might not have noticed but you are in the "Literature" subject. Hover your mouse where it says Literature and find the correct subject by clicking the 'Find More Subjects' button or simply clicking the correct subject in the list that you have already studied. You can also get to the correct subject easily by going to this link: http://www.openstudy.com/study#/groups/Mathematics

OpenStudy (atrineas):

@hjutyu you should post this in the mathematics section you'll get more help there

OpenStudy (savannah_noelle):

let's use the pythogorean theorem; let = length of ramp = l l ² = 6^2*16^2 ....= 36+256 ....= 292 Then find the squareroot of 292 and you get 17.08... Round the answer to 17.1 You should post this in the Mathematics section :)